OpenStudy (anonymous):

find the slope of the graph of the function y=(5x)/(3-x) at (4,20)

OpenStudy (chaise):

Differentiate the equation, and make it equal to zero, solve it and then plug in your values for x and y.

OpenStudy (anonymous):

Really? At (4,20)? Your professor is super mature.

OpenStudy (anonymous):

But here's how to solve it: Using quotient rule y' =[ 5(3-x) - (-1)(5x) ]/(3-x)^2 = [(15-5x) +5x] /[x^2 -6x +9] =[15]/[x^2 -6x +9] Substitute in x=4 and solve.